OpenStudy (anonymous):

PLEASE HELP. Rotate the axes to eliminate the xy term in the equation. then write the equation in standard form... a. x^2-4xy+y^2+9=0 and b. 5x^2-2xy+5y^2-12=0

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

take X=ax+by and Y=bx-ay putting this in the first equation (ax+by)^2-4(ax+by)(bx-ay)+(bx-ay)^2+9=0 this gives u an xy term with co-efficient a^2-b^2,so a^2=b^2 from 2nd equation also u get a^2=b^2....so u can take any pair of a,b....

OpenStudy (anonymous):

Question a and b are separate problems :/ thank you though